a coffee merchant has coffee beans that sell for $9 a per pound and $12 per pound. The two types are to be mixed to create 100 pounds of mixture that will sell for $11.25 per pound. How much of each type of bean should be in used in the mixture?
Answer by checkley71(8403)

12X+9(100-X)=100*11.25
12X+900-9X=1125
3X=1125-900
3X=225
X=225/3
X=75 POUNDS OF $12 COFFEE.
100-75=25 POUNDS OF $9 COFFEE
PROOF
12*75+9*25=1125
900+225=1125
1125=1125